# Readout of multi-level quantum geometry from electronic transport

### v1

@@ Abstract
Introduction.—The quantum geometric tensor (QGT) is a fundamental object in the study of quantum systems and reflects the internal geometry of quantum states in an arbitrary parameter space. It naturally arises in different branches of physics and provides a unified framework for understanding how the quantum states and topology of a system change with changing these parameters [1, 2, 3, 4, 5]. For instance, geometric concepts related with the QGT appear in topological quantum materials [6, 7, 8, 9, 10, 11], affect superfluidity and superconductivity in flat-band systems [12, 13, 14, 15, 16, 17], are relevant for holonomic quantum gates [18, 19, 20, 21, 22] and quantum computing [23, 24, 25, 26, 27, 28, 29, 30], and contribute to nonlinear photoconductivity effects [31, 32, 33, 34]. These concepts have been extended in several directions, including non-Abelian quantum states [35, 36, 37], mixed states [38, 39, 40, 41, 42, 43, 44, 45], and non-Hermitian or open quantum systems [46, 47, 48, 49, 50, 51, 52, 53]. From the experimental side, geometric and topological properties of quantum states have already been measured in a variety of platforms [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66]. However, to this date, the measurement of the full QGT in general electronic transport setups, as sketched in Fig. 1(a), remains an open question.
@@
We consider specific one- and two-parameter drives in the spirit of Ref. [82], which was subsequently applied in both theory [83, 84, 85] and experiment [54, 58, 62] across different platforms. As sketched in Fig. 1, resonantly driving one parameter $\lambda_{\alpha}\to\lambda_{\alpha}+A\cos(\Omega t)$ at a frequency $\Omega$ with amplitude $A$ will give access to the diagonal transition elements $q_{\alpha\alpha,pn}$. A simultaneous drive of two parameters, $(\lambda_{\alpha},\lambda_{\beta})\to(\lambda_{\alpha}+A\cos(\Omega t),\lambda_{\beta}+A\cos(\Omega t-\gamma))$ for $\alpha\neq\beta$, together with specific combinations of conductance measurements taken at different values of the relative phase difference $\gamma$, will give access to the off-diagonal transition elements $q_{\alpha\beta,pn}$.
@@ I Current across a driven center
$G^{\mathrm{adv}}_{ab,\sigma\sigma^{\prime}}(t,t^{\prime})$ $=\frac{i}{\hbar}\theta(t^{\prime}-t)\braket{\{c_{a\sigma}(t),c^{\dagger}_{b\sigma^{\prime}}(t^{\prime})\}}$ (S6b)
@@ IV.1 Example 1: Two-level system
The Hamiltonian $H_{C}$ of the two-level system is defined in the End Matter of the main text in Eq. (13) and the two parameters of the system are $(\lambda_{1},\lambda_{2})=(\theta,\phi)$.
IV.1.1 Drive of $\theta$
Driving $\theta\to\theta+A\cos(\Omega t)$, we have
$W_{k}=\begin{pmatrix}B\cos(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]&Be^{-i\phi}\sin(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]\\ Be^{i\phi}\sin(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]&-B\cos(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]\\ \end{pmatrix}$. (S47)
IV.1.2 Drive of $\phi$
Driving $\phi\to\phi+A\cos(\Omega t)$, we have
@@ IV.1.2 Drive of $\phi$
IV.1.3 Drive of $\theta$ and $\phi$
Driving $\theta\to\theta+A\cos(\Omega t)$ and $\phi\to\phi+A\cos(\Omega t-\gamma)$, we have
@@ IV.2 Example 2: Three-level system
The Hamiltonian $H_{C}$ of the three-level system is defined in Eq. (17) and the three parameters of the system are $(\lambda_{1},\lambda_{2},\lambda_{3})=(\theta,\phi_{1},\phi_{2})$.
IV.2.1 Drive of $\theta$
Driving $\theta\to\theta+A\cos(\Omega t)$, we have
@@ IV.2.1 Drive of $\theta$
IV.2.2 Drive of $\phi_{1}$
Driving $\phi_{1}\to\phi_{1}+A\cos(\Omega t)$, we have
$W_{k}=\begin{pmatrix}0&B\cos(\theta)e^{-i\phi_{1}}[(-i)^{k}J_{k}(A)-\delta_{k0}]&0\\ B\cos(\theta)e^{i\phi_{1}}[i^{k}J_{k}(A)-\delta_{k0}]&0&0\\ 0&0&0\end{pmatrix}$. (S58)
IV.2.3 Drive of $\phi_{2}$
Driving $\phi_{2}\to\phi_{2}+A\cos(\Omega t)$, we have
@@ IV.2.3 Drive of $\phi_{2}$
IV.2.4 Drive of $\theta$ and $\phi_{1}$
Driving $\theta\to\theta+A\cos(\Omega t)$ and $\phi_{1}\to\phi_{1}+A\cos(\Omega t-\gamma)$, we have
@@ IV.2.5 Drive of $\theta$ and $\phi_{2}$
$W_{k}$ $=\left\{\begin{array}[]{ccc}{\tiny\begin{pmatrix}0&Be^{-i\phi_{1}}\cos(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]&0\\ Be^{i\phi_{1}}\cos(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]&0&-iBe^{i\theta}e^{i\phi_{2}}[i^{k}J_{k}(2A)-\delta_{k0}]/2\\ 0&iBe^{-i\theta}e^{-i\phi_{2}}[(-i)^{k}J_{k}(2A)-\delta_{k0}]/2&0\end{pmatrix}}&,&\gamma=0\\ \\ {\tiny\begin{pmatrix}0&Be^{-i\phi_{1}}\cos(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]&0\\ Be^{i\phi_{1}}\cos(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]&0&iBe^{-i\theta}e^{i\phi_{2}}[(-i)^{k}J_{k}(2A)-\delta_{k0}]/2\\ 0&-iBe^{i\theta}e^{-i\phi_{2}}[i^{k}J_{k}(2A)-\delta_{k0}]/2&0\end{pmatrix}}&,&\gamma=\pi\\ \\ {\tiny\begin{pmatrix}0&Be^{-i\phi_{1}}\cos(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]&0\\ Be^{i\phi_{1}}\cos(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]&0&Be^{i\phi_{2}}[\sin(\theta+k\pi/4)J_{k}(\sqrt{2}A)-\delta_{k0}\sin(\theta)]\\ 0&Be^{-i\phi_{2}}[\sin(\theta-k\pi/4)J_{-k}(\sqrt{2}A)-\delta_{k0}\sin(\theta)]&0\end{pmatrix}}&,&\gamma=\frac{\pi}{2}\\ \\ {\tiny\begin{pmatrix}0&Be^{-i\phi_{1}}\cos(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]&0\\ Be^{i\phi_{1}}\cos(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]&0&Be^{i\phi_{2}}[\sin(\theta-k\pi/4)J_{-k}(\sqrt{2}A)-\delta_{k0}\sin(\theta)]\\ 0&Be^{-i\phi_{2}}[\sin(\theta+k\pi/4)J_{k}(\sqrt{2}A)-\delta_{k0}\sin(\theta)]&0\end{pmatrix}}&,&\gamma=-\frac{\pi}{2}\end{array}\right.$. (S75)
IV.2.6 Drive of $\phi_{1}$ and $\phi_{2}$
Driving $\phi_{1}\to\phi_{1}+A\cos(\Omega t)$ and $\phi_{2}\to\phi_{2}+A\cos(\Omega t-\gamma)$, we have

### v2

@@ Abstract
Introduction.—The quantum geometric tensor (QGT) is a fundamental object in the study of quantum systems and reflects the internal geometry of quantum states in an arbitrary parameter space. It naturally arises in different branches of physics and provides a unified framework for understanding how the quantum states and topology of a system change with changing these parameters [1, 2, 3, 4, 5]. For instance, geometric concepts related with the QGT appear in topological quantum materials [6, 7, 8, 9, 10], affect superfluidity and superconductivity in flat-band systems [11, 12, 13, 14, 15, 16], are relevant for holonomic quantum gates [17, 18, 19, 20, 21] and quantum computing [22, 23, 24, 25, 26, 27, 28, 29], and contribute to nonlinear photoconductivity effects [30, 31, 32, 33, 34]. These concepts have been generalized in several directions, including non-Abelian [35, 36, 37, 38] and mixed quantum states [39, 40, 41, 42, 43, 44, 45, 46, 47], non-Hermitian and open quantum systems [48, 49, 50, 51, 52, 53, 54, 55], tensor monopoles [56, 57, 58, 59, 60], and Plücker embeddings [61]. From the experimental side, geometric and topological properties of quantum states have been measured in a variety of platforms [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75]. However, to this date, the measurement of the QGT in general electronic transport setups, sketched in Fig. 1(a), remains an open question.
@@
We consider specific one- and two-parameter drives in the spirit of Ref. [91], which was subsequently applied in both theory [92, 93, 94, 95] and experiment [62, 66, 71] across different platforms. As sketched in Fig. 1, resonantly driving one parameter $\lambda_{\alpha}\to\lambda_{\alpha}+A\cos(\Omega t)$ at a frequency $\Omega$ with amplitude $A$ will give access to the diagonal transition elements $q_{\alpha\alpha,pn}$. A simultaneous drive of two parameters, $(\lambda_{\alpha},\lambda_{\beta})\to(\lambda_{\alpha}+A\cos(\Omega t),\lambda_{\beta}+A\cos(\Omega t-\gamma))$ for $\alpha\neq\beta$, together with specific combinations of conductance measurements taken at different values of the relative phase difference $\gamma$, will give access to the off-diagonal transition elements $q_{\alpha\beta,pn}$.
@@ I Current across a driven center
$G^{\mathrm{adv}}_{ab,\sigma\sigma^{\prime}}(t,t^{\prime})$ $=\frac{i}{\hbar}\theta(t^{\prime}-t)\braket{\{c_{a\sigma}(t),c^{\dagger}_{b\sigma^{\prime}}(t^{\prime})\}},$ (S6b)
@@ IV.1 Example 1: Two-level system
The Hamiltonian $H_{C}$ of the two-level system is defined in the End Matter of the main text in Eq. (13) and the two parameters of the system are $(\lambda_{1},\lambda_{2})=(\theta,\phi)$. Driving only $\theta\to\theta+A\cos(\Omega t)$, we have
$W_{k}=\begin{pmatrix}B\cos(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]&Be^{-i\phi}\sin(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]\\ Be^{i\phi}\sin(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]&-B\cos(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]\\ \end{pmatrix}$, (S47)
while driving only $\phi\to\phi+A\cos(\Omega t)$ yields
@@
The simultaneous drive of $\theta\to\theta+A\cos(\Omega t)$ and $\phi\to\phi+A\cos(\Omega t-\gamma)$ results in
@@ IV.2 Example 2: Three-level system
The Hamiltonian $H_{C}$ of the three-level system is defined in Eq. (17) and the three parameters of the system are $(\lambda_{1},\lambda_{2},\lambda_{3})=(\theta,\phi_{1},\phi_{2})$. Driving only $\theta\to\theta+A\cos(\Omega t)$, we have
@@
Driving only the hopping phases, the drive $\phi_{1}\to\phi_{1}+A\cos(\Omega t)$ yields
$W_{k}=\begin{pmatrix}0&B\cos(\theta)e^{-i\phi_{1}}[(-i)^{k}J_{k}(A)-\delta_{k0}]&0\\ B\cos(\theta)e^{i\phi_{1}}[i^{k}J_{k}(A)-\delta_{k0}]&0&0\\ 0&0&0\end{pmatrix}$, (S58)
while the drive $\phi_{2}\to\phi_{2}+A\cos(\Omega t)$ results in
@@
For driving two parameters simultaneously, we have three possible combinations. Driving $\theta\to\theta+A\cos(\Omega t)$ and $\phi_{1}\to\phi_{1}+A\cos(\Omega t-\gamma)$ gives
@@
$W_{k}$ $=\left\{\begin{array}[]{ccc}{\tiny\begin{pmatrix}0&Be^{-i\phi_{1}}\cos(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]&0\\ Be^{i\phi_{1}}\cos(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]&0&-iBe^{i\theta}e^{i\phi_{2}}[i^{k}J_{k}(2A)-\delta_{k0}]/2\\ 0&iBe^{-i\theta}e^{-i\phi_{2}}[(-i)^{k}J_{k}(2A)-\delta_{k0}]/2&0\end{pmatrix}}&,&\gamma=0\\ \\ {\tiny\begin{pmatrix}0&Be^{-i\phi_{1}}\cos(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]&0\\ Be^{i\phi_{1}}\cos(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]&0&iBe^{-i\theta}e^{i\phi_{2}}[(-i)^{k}J_{k}(2A)-\delta_{k0}]/2\\ 0&-iBe^{i\theta}e^{-i\phi_{2}}[i^{k}J_{k}(2A)-\delta_{k0}]/2&0\end{pmatrix}}&,&\gamma=\pi\\ \\ {\tiny\begin{pmatrix}0&Be^{-i\phi_{1}}\cos(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]&0\\ Be^{i\phi_{1}}\cos(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]&0&Be^{i\phi_{2}}[\sin(\theta+k\pi/4)J_{k}(\sqrt{2}A)-\delta_{k0}\sin(\theta)]\\ 0&Be^{-i\phi_{2}}[\sin(\theta-k\pi/4)J_{-k}(\sqrt{2}A)-\delta_{k0}\sin(\theta)]&0\end{pmatrix}}&,&\gamma=\frac{\pi}{2}\\ \\ {\tiny\begin{pmatrix}0&Be^{-i\phi_{1}}\cos(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]&0\\ Be^{i\phi_{1}}\cos(\theta+k\pi/2)[J_{k}(A)-\delta_{k0}]&0&Be^{i\phi_{2}}[\sin(\theta-k\pi/4)J_{-k}(\sqrt{2}A)-\delta_{k0}\sin(\theta)]\\ 0&Be^{-i\phi_{2}}[\sin(\theta+k\pi/4)J_{k}(\sqrt{2}A)-\delta_{k0}\sin(\theta)]&0\end{pmatrix}}&,&\gamma=-\frac{\pi}{2}\end{array}\right$., (S75)
while driving $\phi_{1}\to\phi_{1}+A\cos(\Omega t)$ and $\phi_{2}\to\phi_{2}+A\cos(\Omega t-\gamma)$, we get